OpenStudy (anonymous):

A helium-neon laser (λ = 738 nm) is used to calibrate a diffraction grating. If the first-order maximum occurs at 22.8°, what is the line spacing?

OpenStudy (anonymous):

d *sin(θ) = m*λ d = m*λ/sin(θ) d =1*7.38 x 10^-7/sin(22.8) d =1.904439527e-6

OpenStudy (jamesj):

Looks about right!